OpenStudy (anonymous):

What is the slope of a line perpendicular to the line whose equation is y = 2x + 5? slope = -1 slope = -1/2 slope = -2

OpenStudy (anonymous):

perpendicular slope= -1/m=-1/2

OpenStudy (anonymous):

the perpendicular slope always equals -1 if the two slopes are multiplied together. and -1/2 times 2 is -1 so that is the perpendicular slope

OpenStudy (anonymous):

does that make sense? @Maddiem

OpenStudy (anonymous):

Yes thanks @GretaKnows

OpenStudy (anonymous):

no problem! :) medal?

OpenStudy (anonymous):

im new so i dont know how to give u one.

OpenStudy (anonymous):

I got one! you just click on best response on the comment that helped you the most.

OpenStudy (anonymous):

Oh cool thanks!